What is the density of Carbon Dioxide gas at


a. Standard temperature and pressure?

 b. 20 degree Celsius and 735 mm Hg?


Note: density = mass/Volume       n = mass/molar mass  


1
Expert's answer
2021-02-01T03:42:55-0500

a. Mass of co2 is equal to the mass of 1 molecule of carbon+mass of 2 molecules of oxygen

Mass=12=(2x16)=44.

N is the mass of 1 molecule which is equal to 44.

Density=mass/volume

Find volume of c02 is equal to v=nRT/P

R=8.314772(molar gas constant)

N=1, T=273k, p=1 or 101325pa

v=(1x8.314772X273)/101325=0.02240249

d=m/v(mass in kg)

d=0.044/0.02240249=1.9784

Density of 1 mole of carbon dioxide at standard temperature and pressure is 1.9784


b. V={8.314772x(20+273)}/(735x133.322)

2436.2282/97991.67

v=0.0248618

d=m/v

=0.044/0.0248618

density=1.7698.